OpenStudy (anonymous):

What is the equation of the ellipse with co-vertices (0, 2), (0, -2) and vertices (3, 0), (-3, 0)? Medal!!

OpenStudy (amistre64):

what is your general equation to fill in?

OpenStudy (anonymous):

I don't know this is why i need help

OpenStudy (amistre64):

the general formula for an ellipse should already be presented to you in your material ...

OpenStudy (michele_laino):

the general equation of an ellipse, is: \[\frac{{{x^2}}}{{{a^2}}} + \frac{{{y^2}}}{{{b^2}}} = 1\] where, in your case, \(a=3,b=2\)

OpenStudy (michele_laino):

please, replace \(a\) with \(3\) and \(b\) with \(2\)